OpenStudy (anonymous):

Piecewise Funtion: A local store owner pays her employees time-and-a-half for overtime. That means if an employee works more than 40 hours per week, the store will pay him 1.5 times his regular hourly wage of $6.50. Write and graph a piecewise function that illustrates an employee's weekly pay, W, as a function of the number of hours, h, that employee works. I thought I figured it out but I'm confused and I really need help to graph it too :( Thank you

OpenStudy (smileyxl3):

Lol, I was asking the same question earlier. I still don't have the answer but you can look at these two to see if it will help http://openstudy.com/updates/516b671ce4b02ec89c5aac8b http://openstudy.com/updates/51073ccee4b069b605a24ab7

OpenStudy (anonymous):

@smileyxl3 thank you!

OpenStudy (anonymous):

The wage is hourly wage up to 40hr, so that goes as y = 6.50 x then from 40 hours up use y = 6.50(40) + (1.5)(6.50)(x-40)

OpenStudy (anonymous):

I'm confused /: so whats my piecewise function

OpenStudy (anonymous):

y = 6.50 x from x = 0 to 40 first piece y = 6.50(40) - (1.5)(6.50)(x -40 from x = 40 and more, 2nd piece